OpenStudy (anonymous):

how do i solve x^2-16

myininaya (myininaya):

there is nothing to solve

OpenStudy (anonymous):

x^2-16 factors into (x-4)(x+4) = 0 (x-4) = 0 (x+4) = 0 x= 4 and -4